# Exploiting Student Parallelism for Low-latency GPU Inference of BERT-like Models in Online Services

### v1

@@ 4.3.2. Accuracy Analysis
As shown in Table 4, we first conduct accuracy ablation experiments to verify the effectiveness of our training methods to maintain accuracy on the MRPC dataset, as stated in Table 3. When we directly use all trained students without adaptive student pruning, we observe that F1 score decreases a little from 86.9 to 86.8. After removing the stacking distillation for better refinement, the F1 score further drops to 86.0, which is slightly lower than the TinyBERT-4L312D. We further replace the boosting ensemble with a simple bagging ensemble of multiple small students, and the performance will further decrease. The same trend also happens on the other datasets.
@@ 4.3.3. Latency Analysis
We also make breaking-down experiments to evaluate the influences and contributions of different design parts on inference latency. The results show that latency increases dramatically when we increase the student’s model depth to 4 layers, which indicates that reducing serial computation is the key to low latency. After we pad all data samples into the maximum length, the average latency increase to be similar to tail latency. It means that the variances of Student Parallelism latency mainly come from the irregular sentence lengths. We further employ the waiting queue to batch enough samples for data parallelism, then average latency increase a lot and tail latency even get close to the overall time budget.
@@
5. Conclusion
We propose Student Parallelism to adopt student parallelism by distilling the deep BERT-like model into an equivalent group of parallel shallow students. Every student model can have decoupled and homogenous architecture of fewer layers for low inference latency, while all student models work together to maintain accuracy via boosting ensemble and stacking distillation. And adaptive student pruning can temporally drop some students to improve the throughput for workload burst. We further make specialized system designs for student parallelism to minimize waiting and padding. Comprehensive experiments have verified the effectiveness of our work in terms of latency, throughput, and accuracy.
Appendix A Convergence Analysis
@@ Appendix B Implementation Details
B.1. Testbed and Environment
@@
B.2. Offline Training Implementation
In the offline training, we consider all 24 compact BERT variants [39] and all DEBERTA variants as the student candidates to find the lowest one for different teachers. In practice, we use the BERT-2L256D [39] for BERT-base, BERT-6L768D for BERT-large, and Deberta-12L768D for DEBERTA-xlarge as the backbone network and initialization of student models respectively. Following previous works [16, 43], we conduct the same data augmentation on the down-stream datasets. We conduct a grid search on various training hyperparameters to find the best for every dataset, including epoch number, batch size, learning rate, $\lambda$ weights, and data sampling ratio. And in practice, we optimize the $\alpha_{i}$ together with the student, since they are both differentiable.
B.3. Online Inference Implementation
For online inference, the global dispatcher to assign samples among different nodes in the round-robin method. And every node employs message queues to dispatch samples and collect the results in the master-worker architecture. The master process is responsible for logical controlling, including resource management, the length-aware buffer, and task dispatching. And workers host all concurrent student groups on all the GPUs. Once a user submits an inference request, the master preprocesses, buffers, and then dispatches the inference samples. And any idle worker can fetch the inference task from the message queue and return the results. We employ the Multi-Process Service (MPS) [28] to make multiple students of different groups share the same GPUs without interference.
B.3.1. Hyperparameters for Student Allocation
@@
B.3.2. Dispatch with Length-Aware Buffer
Figure 7. Length-aware Buffer: New sequence sample (e.g., the red one whose length is 30) can be merged with the previous element into a small batch, since they share the same bin and the element size is not full. Otherwise, the new sequence sample (e.g., the blue one whose length is 120) is added to the buffer tail as a new element.

### v2

@@ 4.3.2. Accuracy Analysis
As shown in Table 4, we first conduct accuracy ablation experiments to verify the effectiveness of our training methods to maintain accuracy on the MRPC dataset, as stated in Table 3. When we directly use all trained students without adaptive student pruning, we observe that the F1 score decreases a little from 86.9 to 86.8. After removing the stacking distillation for better refinement, the F1 score further drops to 86.0, which is slightly lower than the TinyBERT-4L312D. We further replace the boosting ensemble with a simple bagging ensemble of multiple small students, and the performance will further decrease. The same trend also happens on the other datasets.
@@ 4.3.3. Latency Analysis
We also make breaking-down experiments to evaluate the influences and contributions of different design parts on inference latency. The results show that latency increases dramatically when we increase the student’s model depth to 4 layers, which indicates that reducing serial computation is the key to low latency. After we pad all data samples into the maximum length, the average latency increases to be similar to tail latency. It means that the variances of Student Parallelism latency mainly come from the irregular sentence lengths. We further employ the waiting queue to batch enough samples for data parallelism, then average latency increases a lot and tail latency even gets close to the overall time budget.
@@
5. Discussion
Extention to Generative Models. The generative models generate tokens step by step, and every step makes a classification decision. Therefore, our work can be naturally extended to speed up every token classification. However, we do not conduct such experiments due to three main reasons: (1) There is no golden standard (e.g., an objective metric like accuracy or F1 in discriminative tasks) to measure how well a sentence is generated (perplexity and BLEU are arguably only the indirect references instead of the golden one). It is relatively harder to verify if it gets good prediction quality. (2) We didn’t find the 2-layer pre-trained GPT like BERT-2L from well-read students (turc2019well, ) to initialize the generative students. And we cannot afford to collect such large datasets and conduct large-scale pre-training tasks. (3) Most importantly, we believe that speeding up BERT is non-trivial. They have been widely adopted by online services related to text mining and information retrieval. Information retrieval plays important roles like RAG (lewis2020retrieval, ) in the future because GPT suffers from the risks of hallucination.
Training Cost. Student Parallelism has a similar time cost to other model compressions(e.g., 1.35x the training time of DynaBert with comparable accuracy), when using one GPU for training. Although we need to train multiple students, the student is a much smaller (e.g., 25% depth and 50% width) pre-trained model compared with other baselines. Each student is only trained on the sub-sampled dataset(e.g., 20%) based on the residual error in the boost ensemble learning. Pruning and quantization also need extra training or calibration time. Furthermore, the training of multiple students can be easily accelerated by multiple GPUs, without tuning parameters for larger batch sizes. Specifically, we can use one GPU to train the current student and distribute the inference of previous students on different GPUs. Most importantly, online services like search engines can have much more inference samples (e.g., billions a day) than training samples, which can amortize the training costs.
6. Conclusion and Future Work
We propose Student Parallelism to adopt student parallelism to trade more students for less model depth. Every student model can have a decoupled and homogeneous architecture of fewer layers for low inference latency because all student models work together to maintain accuracy via boosting ensemble and stacking distillation. And adaptive student pruning can temporally drop some students to improve the throughput for workload burst. We further make specialized designs for student parallelism to minimize waiting and padding. Comprehensive experiments with real online workloads have verified the effectiveness of our work in terms of latency, throughput, and accuracy. In future work, we will extend it to generative models like GPT and reinforcement learning.
Appendix A Theory Analysis
In this section, we first conduct the convergence analysis on our training methods based on boosting ensemble, which shows the residual training loss converges to the infimum as more students added during the training. Then we make the quantitative analysis on the inference efficiency by comparing our proposed Student Parallelism and other typical baselines.
A.1. Convergence Analysis
@@
A.2. Efficiency Analysis for GPU Inference
To make the quantitative analysis and comparison, we first build a performance model for the inference latency and throughput. In this model, we consider all model and input factors, including model depth $D$, model width $W$, batch size $B$, and input length $N$. Moreover, we also consider the time cost of the waiting queue $Q(B)$ that collect the maximum $B$ samples, and the parallel computation capability $C$, the PCI-E transferring bandwidth $T$ from the host to the GPU, the GPU number $G$, and the total parallel model or student group number $M$ on the cluster. Therefore, we can model the inference latency of BERT-like models as follows:
(14) $\text{Latency}=\Bigl{\lceil}\frac{WBN^{2}M}{CG}\Bigr{\rceil}+Q(B)+\frac{BN}{T}\$
The first term is the inference computation time on the GPU, in which $\Bigl{\lceil}\frac{WBN^{2}M}{CG}\Bigr{\rceil}$ is computation time of one layer and the ceiling symbol means it is constant if there is enough parallel computation capability $C$. And the second term $Q(B)$ stands for the waiting time of collecting the data batch, which depends on the workload and batch size B. The last term means the PCI-E transferring time of the data size $BM$. Furthermore, we can further model the throughput per GPU based on the inference latency as follows:
(15) Throughput Per $\text{Throughput Per GPU}=\frac{BM}{\text{Latency}\cdot G}$
Obviously, the total sample number $BM$, the GPU number $G$, and the latency determine the throughput per GPU together.
Depth Width Batch Size Length M Q
BERT-base 12 768 $\sim$10 128 G Yes
TinyBERT 4 312 $\sim$10 128 G Yes
DynaBERT 6 192 $\sim$10 128 G Yes
DeeBERT 1-12 768 $\sim$10 128 G Yes
Cocktail $\max(D_{i})$ $\sum_{i}W_{i}$ $\sim$10 128 G Yes
Student Parallelism 2 (1-3)*256 $\leq 4$ 8(n%8+1) 4G None
Table 5. The comparison on all related factors with latency and throughput: M is the number of running model or student group, G is the GPU number, and Q is the waiting queue.
To analyze our advantages in efficiency, we then make quantitative comparisons between Student Parallelism and some representative baselines in optimizing the BERT-base on the MRPC dataset. As all the related factors shown in Table 5, Student Parallelism can outperform the other baselines due to its model architecture and system designs. Specifically, Student Parallelism can achieve the lowest model depth of only 2 layers to significantly decreases the latency. Although the overall model width of 3 students brings in more parallel operators, the GPUs have enough parallel computation capability to speed them up. Because Student Parallelism realizes the direct inference on different individual samples, there is no waiting $Q$ and large batch $B$ for PCI-E transferring. And the single sample and small batches of 2$\sim$4 samples in the similar length has the input length as the closest bin length 8(n%8+1). In terms of throughput, the multiple students sharing the same GPU make the $M$ can be several times larger than $G$, leading to large enough total concurrent sample number $BM$ to fully utilize the GPUs. Additionally, adaptive student pruning can reduce the student number as low as only 1, which can improve the student group number $M$ to achieve the highest thought.
In contrast, the other baselines suffer from the model architecture unfriendly to GPUs and extra costs for data parallelism. The knowledge distillation baseline TinyBERT has 4 layers fewer than any other baselines, but it is 2 times larger than Student Parallelism. In adaptive computing baseline DeeBERT, some ”easy” sample can exit from the bottom layers, while other ”hard” ones still have to go throughpu all 12 layers. Because Cocktail is the bagging ensemble of different models, its model depth is determined by the straggler that have the largest depth max($D_{i}$). Although the pruning baseline DynaBERT have the smallest model width of 192 hidden dimensions in all methods, it mainly reduces the parallel operators that can be efficiently accelerated by GPUs. And all baselines need the large enough data batches for data parallelism, brining in considerable waiting queue costs $Q$ and large PCI-E transferring time for the whole batch. Traditionally, they have to pad all short samples to have the same maximum length $N$ as 128. Otherwise, they need to suffer from extra compiling cost or computation divergence for the ragged batch. Based on data parallelism, all the baselines only run one model on every GPU to process different data batches respectively, resulting in $M=G$.
@@ Appendix B Implementation Details
B.1. Additonal Experiment Settings
By default in Fig. 1, the model is a 12-layer BERT-base with 768 hidden dimensions, and the inference data batch consists of 8 samples with length of 256. In every sub-figure, we choose one as the changing variable and fix the others to measure the latency and throughput and show the influences of the changing variable.
B.2. Testbed and Environment
@@
B.3. Training Implementation Details
In the offline training, we consider all 24 compact BERT variants [44] and all DEBERTA variants as the student candidates to find the lowest one for different teachers. In practice, we use the BERT-2L256D [44] for BERT-base, BERT-6L768D for BERT-large, and Deberta-12L768D for DEBERTA-xlarge as the backbone network and initialization of student models respectively. Following previous works [19, 49], we conduct the same data augmentation on the down-stream datasets. We conduct a grid search on various training hyperparameters to find the best for every dataset, including epoch number, batch size, learning rate, $\lambda$ weights, and data sampling ratio. And in practice, we optimize the $\alpha_{i}$ together with the student, since they are both differentiable. The training epoch number is early stopped by the validation loss. lr=5e-5 and betas=(0.9, 0.999) for Adam optimizer, $\lambda=1$ works fine in most cases, data sampling ratio is usually 0.2 (smaller in some large datasets).
The student number is determined by the validation loss. When conducting student distillation, we continue to add new students until the validation loss does not decrease any more. Then we conduct adaptive student pruning and measure the validation loss to decide the number of students to maintain. To determine the student model depth, we conduct the binary search from one layer to 25% layers of the teacher model. We choose the smallest model depth that can reach the target accuracy threshold (i.e., 95% of the teacher’s accuracy)
B.4. Online Inference Implementation
For online inference, the global dispatcher to assign samples among different nodes in the round-robin method. And every node employs message queues to dispatch samples and collect the results in the master-worker architecture. The master process is responsible for logical controlling, including resource management, the length-aware buffer, and task dispatching. And workers host all concurrent student groups on all the GPUs. Once a user submits an inference request, the master preprocesses, buffers, and then dispatches the inference samples. And any idle worker can fetch the inference task from the message queue and return the results. We employ the Multi-Process Service (MPS) [33] to make multiple students of different groups share the same GPUs without interference. As shown in Fig 4  sys has the following procedures during the online inference phase:
  1. (1)
  2. (2)
    User submits a request triggering one or more inference samples.
  3. (3)
  4. (4)
  5. (5)
  6. (6)
  7. (7)
    It returns results to users.
B.4.1. Details about Student Allocation
Student Parallelism employs NVIDIA Multi-Process Service (MPS) [33] to make multiple students from different groups share the same GPU without interference. It can guarantee different students on the same GPU have separate computation units and memory bandwidth to run concurrently. Since different individual samples do not arrive at the same time, different student groups can concurrently process them by overlapping between data transferring to GPU and inference computation as shown in Figure 7.
@@ B.4.1. Details about Student Allocation
B.4.2. Details about Sample Dispatching based on Length-aware Buffer
@@
B.4.3. Student Number Adaptation
If there are intensive workloads to fulfill the length-aware buffer, Student Parallelism can further decrease the student number (§ 3.2) to run more student groups. According to the workload intensity and resource usage, the resource manager dynamically adjust the student number to host different numbers of student groups on the available GPUs. It keeps tracking the buffer size and the idle group number as indicators of workload intensity and resource usage. If the buffer size reaches the threshold (i.e. the total student group number), Student Parallelism drops one last-trained student models to start more groups, so that the waiting time of buffered samples will not exceed the inference time. And if the length-aware buffer is empty for a certain time window (e.g., 2 minutes), Student Parallelism checks whether total student number of all idle groups is larger than occupied group number. If so, it means there are enough free resources to increase the student number by one for all the occupied student groups. Then Student Parallelism can make better use of computation resources of the idle students to improve accuracy. To avoid cold starting for the entire environment, the inactivated student instances of different student number settings can be preloaded in the host memory to realize hot standby.